# Test-time Computing: from System-1 Thinking to System-2 Thinking

### v1

@@ 6 Conclusion
In this paper, we conduct a comprehensive survey of existing works on test-time computing. We introduce various test-time computing methods in System-1 and System-2 models, and look forward to future directions for this field. We believe test-time computing can help models handle complex real-world distributions and tasks better, making it a promising path for advancing LLMs toward cognitive intelligence. We hope this paper will promote further research in this area.
@@
Test-time computing, especially the strategies in System-2, is evolving rapidly. While we have made efforts to provide a comprehensive survey of existing research, it is challenging to cover all the latest developments. This review includes papers up to November 2024, with more recent advancements to be updated in future versions. TTA has seen many successful applications and task-specific strategies in CV tasks. Since the primary audience of our paper is researchers in NLP, we do not systematically present these works, and interested readers can refer to Liang et al. (2024a) for details.

### v2

@@ 7 Conclusion
In this paper, we conduct a comprehensive survey of existing works on test-time compute. We introduce various test-time compute methods in System-1 and System-2 models, and look forward to future directions for this field. We believe test-time compute can help models handle complex real-world distributions and tasks better, making it a promising path for advancing LLMs toward cognitive intelligence. We hope this paper will promote further research in this area.
@@
Test-time compute, especially the strategies in System-2, is evolving rapidly. While we have made efforts to provide a comprehensive survey of existing research, it is challenging to cover all the latest developments. This review includes papers up to January 2025, with more recent advancements to be updated in future versions. TTA has seen many successful applications and task-specific strategies in CV tasks. Since the primary audience of our paper is researchers in NLP, we do not systematically present these works, and interested readers can refer to Liang et al. (2024a) for details.